OpenStudy (anonymous):

write in lowest terms as a whole or mixed number 8/9 x 72

OpenStudy (anonymous):

Is the equation \[\frac{8}{9}\times72\]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

that = 64 but im not sure what the problem is asking

OpenStudy (anonymous):

write in lowest terms thats all

OpenStudy (anonymous):

The question is asking HOW: \[\frac{8}{9}\times72=8\times\frac{72}{9}=8\times8=64\]

OpenStudy (anonymous):

thanks you two :)

OpenStudy (anonymous):

No problem!